Bonanza Co. manufactures products X and Y from a joint process that also yields a by-product, Z. Revenue from sales of Z is treated as a reduction of joint costs. Additional information is as follows:

PRODUCTS
X Y Z TOTAL
Units produced 34,000 34,000 17,000 85,000
Joint costs ? ? ? $ 332,000
Sales value at split-off $ 510,000 $ 255,000 $ 17,000 $ 782,000


Joint costs were allocated using the net realizable value method at the split-off point. The joint costs allocated to product X were

Multiple Choice

  • $255,000.

  • $127,500.

  • $142,800.

  • $210,000.
Answer #1

Joint cost allocable = 332000-17000 = 315000

Joint Cost allocation to product X = Total Joint Cost/Total net realizable value*Net realizable value of product X

= 315000/765000*510000

Joint cost allocation to product X = 210000

So answer is d) $210000